OpenStudy (anonymous):

How fast must a ball be rolled along the surface of a 70-cm high table so that when it rolls off the edge it will strike the floor at the same distance (70cm) from the point directly below the edge of the table?

OpenStudy (amistre64):

change centimeters to meters and use -4.9 as the force of gravity to see how much time it takes to fall 70 cm

OpenStudy (amistre64):

then you wil know that distance and the time it takes and that is a ratio of speed

OpenStudy (amistre64):

-4.9t^2 + h = 0 when t^2 = h/4.9; so t = sqrt(h/4.9) \[\frac{h}{\sqrt{h}/\sqrt{4.9}}\] \[\frac{h\sqrt{4.9}}{\sqrt{h}}\] \[\sqrt{4.9h}\] maybe?

OpenStudy (amistre64):

70 (.01 m = 1 cm) = h .7 m = 70 cm = h \[\sqrt{4.9(.7)}\ mps\] \[\sqrt{3.43}\ mps\]

OpenStudy (amistre64):

i believe i might have a simple error in there someplace

OpenStudy (amistre64):

nah, its right, it just looks odd when i do it another way

OpenStudy (anonymous):

horizontal movement: s = distance travelled = 0.70 m let v = velocity as it leaves table t = time to reasch floor v = 0.70 / t vertical movement s = 0.70 u = initial vertical velocoty = 0 t = time taken g = 9.8 m s-2 0.70 = 0t + 0.5*9.8 * t^2 t^2 = 0.70 / 4.9 = 1/7 t = 1 / sqrt7 plug this into equation for horizontal movement to get v

OpenStudy (anonymous):

i'm stretching my memory of mechanics but i believe this is correct

OpenStudy (anonymous):

lets see what amis comes up with

OpenStudy (amistre64):

its correct

OpenStudy (amistre64):

\[\frac{.7}{1/\sqrt{7}}=.7\sqrt{7}=\sqrt{3.43}=etc...\]